OpenStudy (anonymous):

what is the equation of the line with an x-intercept of -1 and a y-intercept of 2

OpenStudy (anonymous):

okay so can you write the x intercept and y intercept as coordinates

OpenStudy (mathmath333):



OpenStudy (mathmath333):

when x=-1,y=0 when y=2,x=0

OpenStudy (anonymous):

If it is like mathmath333 you can take the slope from that, and use this equation: y-y1=m(x-x1)

OpenStudy (amistre64):

\[ax + by = c\] \[\frac acx + \frac bcy = 1\] \[\frac x{c/a} + \frac y{c/b} = 1\] when x=0, y = c/b when y=0, x = c/a